OpenStudy (anonymous):

x+1/x = 2 x^100 + 1/x^100 = ? x+1/x = 2 x^100 + 1/x^100 = ? @Mathematics

hero (hero):

x = 1

OpenStudy (anonymous):

Solution please..

hero (hero):

Well, it depends I think it's 2

OpenStudy (anonymous):

yes, it's 2

hero (hero):

Good

OpenStudy (anonymous):

what cat?

OpenStudy (anonymous):

Can anyone explain the solution.

OpenStudy (anonymous):

It's not possible, it's in addition: x+1/x = 2 x+1 = 2x this is wrong It should be : x^2+1 = 2x

OpenStudy (anonymous):

x+1\[\neq\] 2x here

hero (hero):

(x+1)/x = 2 or x + (1/x) = 2 Which one is it?

OpenStudy (anonymous):

First we have to make the denominator equal.

hero (hero):

Maybe you should think about using parentheses when you post problems. I'm sick of helping students who don't know how to properly use parentheses when posting fractions linearly

hero (hero):

If you don't use parentheses, then you haven't written the fraction properly

OpenStudy (anonymous):

Oh, it's my mistake. You gave correct answer, I thought you understood the question.

hero (hero):

abdul, post the original problem again and use parentheses

OpenStudy (anonymous):

\[x+\left( 1/x \right)=2\] \[x ^{100}+\left( 1/x ^{100} \right)=?\]

hero (hero):

This is exactly why I hate the setup they have here

hero (hero):

x+ (1/x) = 2 (x^2/x) + (1/x) = 2 (x^2+1)/x = 2 x^2+1 = 2x x^2-2x+ 1 = 0 x^2-x-x+1 = 0 x(x-1)-1(x-1) = 0 (x-1)(x-1) = 0 (x-1)^2=0 x-1 =0 x =1

hero (hero):

x^100 + (1/x^100) = ? 1^100 + (1/1^100) = ? 1 + 1/1 = ? 1+1 = 2

hero (hero):

Is that good enough?

hero (hero):

Either way, whether you do it right or wrong, you still get 2

OpenStudy (anonymous):

When 2x comes to left side will it not get divided?

hero (hero):

Nope. You never divide by the variable you're solving for

hero (hero):

Remember that forever

OpenStudy (anonymous):

Thanks Hero, Your name best suits you..

hero (hero):

I guess

OpenStudy (anonymous):

or may be it should be SUPER HERO

hero (hero):

In this case, it wouldn't matter, but in other cases, dividing by the variable you're solving for would eliminate a solution set and therefore, once you solved for the variable, you wouldn't have the full solution, only part of it. And thus, a partial solution is not a full solution and you would either get points taken off or it be marked wrong altogether.

OpenStudy (anonymous):

Yeah, I understand

hero (hero):

If you run into problems similar to this, know that your job is to rescue the variables from the denominator and then keep them out of the denominator.

hero (hero):

So you're the super hero, and you must save the variables so that they may have all of their solutions and live happily ever after.

OpenStudy (anonymous):

Then you are the Great Super Hero, because you only taught to save the variables.

hero (hero):

:D

OpenStudy (anonymous):

Which place are you from?

hero (hero):

I'm from a place far far away.

OpenStudy (anonymous):

country?

hero (hero):

Think Galaxy, not country

OpenStudy (anonymous):

Fine, which Special Galaxy other than Milky Way

hero (hero):

Pegasus Galaxy, the other side of the known universe. Me and other Great Super Heroes spend our lives saving variables

OpenStudy (anonymous):

It would take radio signals million of years to reach earth..

OpenStudy (anonymous):

but your taking just a second to reply?

hero (hero):

I came here in a spaceship that travels faster than the speed of light. (Yes such vehicles exist)